OpenStudy (anonymous):

Three 20 cm-long rods form an equilateral triangle. Two of the rods are charged to + 14 nC, the third to - 14nC.What is the electric field strength at the center of the triangle?

OpenStudy (anonymous):

All i need is help on how to start it.

OpenStudy (turingtest):

Well it seems we're gonna want to start with Gauss' law to find the field due to each rod. I'm assuming that is applicable here, but that is a simplification.

OpenStudy (turingtest):

\[\oint\vec E\cdot d\vec A=\frac q{\epsilon_0}\]\[E(2\pi r\ell)=\frac{\lambda\ell}{\epsilon_0}\to\vec E=\frac{\lambda}{2\pi r}\vec r\]so now that we have our field formula we can get our three vectors at the point and superimpose them.

OpenStudy (turingtest):

by symmetry the situation will be this way:

OpenStudy (turingtest):

the angles should be 60/2=30degrees